Sub:Financial Accounting and analysis
Question:
b. Prepare the cash flow statement from investing activities of Alpha Creative Ltd for the
year ended March31, 2019
Particulars Amount
Plant acquired 160000
Claim received for loss of plant in fire 45500
Unsecured loans given to subsidiaries 595000
Interest on loan received from subsidiary companies 72500
•Also give reasons for the classification of above activities as inflow /outflow
1
Expert's answer
2019-12-02T09:24:06-0500

A Cash Flow statement is a statement showing changes in cash position of the firm from one period to another. It explains the inflows (receipts) and outflows (disbursements) of cash over a period of time. The inflows of cash may occur from sale of goods, sale of assets, receipts from debtors, interest, dividend, rent, issue of new shares and debentures, raising of loans, short-term borrowing, etc. The cash outflows may occur on account of purchase of goods, purchase of assets, payment of loans loss on operations, payment of tax and dividend, etc.


Cash Inflow mean the cash which comes in from any sourse wheter its operating activity, Investing or financing. 

Cash Outflow is opposite of Cash inflow.


Cash receipts: 45500 + 72500 = 118000.

Cash paid: 160000 + 595000 = (755000).

Net cash flow: 118000 - 755000 = (637000).